OpenStudy (anonymous):

0^0 = ? reason too plz...

OpenStudy (anonymous):

it is undefined......a^0=a^(x-x)=a^x/a^x putting a=0, this gives 0/0....so undefined...

OpenStudy (anonymous):

contradicting dipankar study..

OpenStudy (anonymous):

pranav89 is wrong...ik kishan its an indeterminate form

OpenStudy (anonymous):

thnx..

OpenStudy (anonymous):

The subject is controversial but read this link: http://www.faqs.org/faqs/sci-math-faq/specialnumbers/0to0/

OpenStudy (anonymous):

0^0 is undefinedddddddddddddddddd

OpenStudy (anonymous):

0^0 only tends to 1

OpenStudy (anonymous):

pranav89 there is no controversy about this..:) lol.. yah lim(x->0)x^0=1

OpenStudy (anonymous):

Yeah - the first comment was a mistake - did not get my facts right! sorry! ik kishan and saubhik - you were right. 0^0 tends to 1

OpenStudy (anonymous):

\[0^0 =1\] ? news to me